# Neural-Enhanced Born Approximation for Inverse Scattering

### v1

@@ Abstract
Time-harmonic, acoustic inverse scattering concerns the ill-posed and nonlinear problem of determining the refractive index of an inaccessible, penetrable scatterer based on far-field wave scattering data. When the scattering is weak, the Born approximation provides a fast linearized model for recovering the shape and material properties of a scatterer. We develop two neural network algorithms–Born-CNN (BCNN) and CNN-Born (CNNB)–to correct the Born approximation when the scattering is not weak. BCNN applies a post-correction to the Born reconstruction, while CNNB pre-corrects the data. Both methods leverage the Born approximation’s excellent fidelity in weak scattering, while extending its applicability beyond its theoretical limits. CNNB particularly exhibits a strong generalization to noisy and absorbing scatterers. Our approach provides a direct and robust method for obtaining the refractive index, extending the utility of the Born approximation to regimes where the traditional method fails.
@@ 1 Introduction
The time-harmonic inverse scattering problem concerns the determination of the properties, such as shape or material composition, of an inaccessible object from remote measurements of wave scattering data. Problems of this type arise in seismology, medical imaging, and radar applications. Due to the significance of the applications, there has been extensive work on various algorithms for this problem (see for example [5]).
We shall study a particular inverse scattering problem: the inverse medium problem for the Helmholtz equation. In this case, it is desired to reconstruct the refractive index of a bounded penetrable scatterer from far field acoustic data. The major complication is that this inverse scattering problem is both ill-posed and nonlinear [5, Theorem 4.21, page 448]. Current approaches can be divided into two broad classes: 1) quantitative methods and 2) qualitative methods. A quantitative method attempts to reconstruct the scatterer directly. Often this involves a non-linear and regularized optimization problem that seeks a reconstruction corresponding to a far field pattern that matches measurement data. This is computationally intensive and can fail due to local minima [5]. Another quantitative approach is to assume that the scattering is weak and to use the inverse Born approximation. We shall discuss this in more detail shortly since correcting this approach is the subject of our paper. In contrast to these quantitative methods, the goal of a qualitative method such as the Linear Sampling Method [4, 5] is to approximate the support or boundary of the scatterer. Approaches of this type do not involve optimization or the solution of the forward problem. They do not require strong a priori assumptions, but cannot distinguish different materials in the field of view.
@@
The novelty of this paper is to suggest and test two schemes combining the Born approximation with a neural network to extend the domain of applicability of the Born approximation. We compare the resulting predictions to the standard regularized inverse Born approximation, and to a CNN trained simply to invert the data along the lines of [9]. After optimizing the networks and training on simple data generated by a few circular scatterers, we show that the CNNB model is remarkably stable to added noise. Limited testing on more complex, out-of-distribution scatterers reveals that all methods improve the fidelity of the reconstructions for strong scatterers, but the inclusion of the Born approximation in the two proposed schemes produces generally better results. The study suggests that combining CNNs and the Born approximation has promise in solving inverse scattering problems and that the training phase does not need to include close copies of the scatterers.
@@ 2 The Inverse scattering problem
Figure 1: An illustration of the scattering problem in a search space $\Omega$. The incident field $u^{i}$ travels with direction of propagation $\varphi$ and interacts with the compactly supported scatterer $\eta(\textbf{x})$. The scattered field $u^{s}$ propagates outward with direction $\theta$. The grey box is the a priori chosen region where the unknown scatters lie.
@@
$u^{i}(\mathbf{x},\mathbf{d})=\exp(\mathrm{i}\mkern 1.0muk\mathbf{x}\cdot\mathbf{d}),\quad\mathbf{d}=\langle\cos\varphi,\sin\varphi\rangle$
@@
$\triangle u+k^{2}\eta(\mathbf{x})u=0\text{ in }$ $\mathbb{R}^{2},$ (1)
@@
Under the conditions given above, equations (1)-(3) have a unique solution for any $k>0$ [5]. The Forward Problem consists of solving the above linear well-posed problem given $\mathbf{d}$, $k$ and $\eta$.
It follows from the fact that $u^{s}$ satisfies the Helmholtz equation and the radiation condition that $u^{s}$ exhibits an asymptotic expansion as an outgoing cylindrical wave for $|x|$ sufficiently large:
@@
The Inverse Problem that we wish to solve is to determine $\mu(\textbf{x})$ (equivalently, $\eta(\mathbf{x})$, the square of the refractive index) given the far field pattern $u_{\infty}(\hat{\mathbf{x}},\mathbf{d})$ for all $\hat{\mathbf{x}}$ and d on $\mathbb{S}^{1}$ (in practice, only a finite set of $\hat{\mathbf{x}}$ and d are used). Here we assume data is given for a single fixed wave number $k>0$. This problem is non-linear and ill-posed [5].
In this study, we use synthetic scattering data generated through a standard finite element approach to approximating (1)-(3). For each incident direction d, the total field $u$ in a neighborhood of the scatterer is computed using the Netgen package [20] with 4th-order elements and a mesh-size request of one-eighth of the local wavelength of the wave. The Sommerfeld radiation condition is handled through a radial Perfectly Matched Layer (PML) implemented by Netgen using complex stretching as discussed in [3]. To generate an approximate far field pattern, we follow [17] to map the near field to a far field pattern.
@@ 2.1 The Born approximation
$H^{1}_{loc}(\mathbb{R}^{2})=\{u:\mathbb{R}^{2}\to\mathbb{C}\>|\>u|_{B_{r}}\in H^{1}(B_{r})\>\>\forall\>\>r>0\},$ (5)
where $B_{r}=B(r,0)$ denotes a ball of radius $r$ centered at the origin. In particular, let $\Omega=B_{R}$ be the smallest such ball containing the support of $\mu$. It can be shown that if $u\in H^{1}_{loc}(\mathbb{R}^{2})$ is a solution to the scattering problem (1)-(3), then $u|_{\Omega}\in L^{2}(\Omega)$ and satisfies the Lippmann-Schwinger equation:
@@
provided that $\|T\|_{\infty}<1$. A sufficient condition for this is [18]
@@
where ${\mathscr{B}}^{*}$ is the $L^{2}$-adjoint of ${\mathscr{B}}$, and $\gamma>0$ is a fixed regularization parameter determined a priori. Note that fast methods exist to compute this approximation (including the NN approach of Zhou [25] or the low rank approximation method of [24]) but we do not use them here.
@@ 3.1 Born discretization
$(U_{\infty})_{i,j}\approx\frac{k^{2}}{\sqrt{8\pi k}}\exp\left(\frac{\mathrm{i}\mkern 1.0mu\pi}{4}\right)h^{2}\sum_{p,q=1}^{n}\exp(ik(\textbf{d}_{j}-\hat{\textbf{x}}_{i})\cdot\textbf{y}_{p,q})M_{p,q},$ (12)
@@
where $\mathscr{B}_{\rm disc}$ encodes the (2+2)-discretization in both angular dimensions and both spatial dimensions, and $\tau_{\mu}$ is the unknown error for the Born approximation. To utilize linear algebraic methods, we can use standard tensor unfolding methods to collapse $\mathscr{B}_{\rm disc}$ along the spatial and angular dimensions to produce the aforementioned stacked Born operator $\mathcal{B}\in\mathbb{C}^{N_{R}N_{S}\times N_{G}^{2}}$. We thus arrive at the following formulation:
@@
where $\vec{u}_{\infty}=\operatorname{vec}(U_{\infty})$ and $\vec{\mu}=\operatorname{vec}(M)$ are the vectorized formats (through appropriate reshaping) of the far field and scatterer matrices respectively. The discrete inverse Born approximation then predict the nodal values of $\mu$ by ignoring the error term $\vec{\tau}_{\mu}$ and using
@@ 3.2 Data generation
Figure 2: Two realizations of the random training data. Each row displays a density plot of $\eta$ for a sample scatterer (left panel) alongside the real and imaginary components of its computed far-field pattern. The far field pattern is the data for the inverse scattering problem.
For training, we want to use simple scatterers. Here we create the scatterer field $\eta(\textbf{x})$ as the union of $N_{c}$ piecewise-constant circles overlaying a homogeneous background of air, where we uniformly sample $N_{c}\in\{1,2,3\}$. The $i$th circle is assigned a constant value $\eta_{i}$ sampled uniformly from $\mathcal{U}(1.1,1.8)$, with the background air set to $\eta=1$ (we shall see that $\eta\approx 1.8$ is outside the weak scattering regime). The radius and position of each circle are sampled uniformly from $\mathcal{U}(0.1,0.3)$ and $\mathcal{U}(-0.7,0.7)$ respectively. In selecting the training data, for simplicity of mesh generation, we enforce that the circles cannot overlap and must be fully contained in the search domain $[-1,1]^{2}$. We generate 20,000 samples for training and validation (under an 80-20% split) with an additional 4,000 samples for testing purposes. A visualization of a typical training sample can be seen in Figure 2.
The validity of the Born approximation is contingent upon weak scattering of the incident field. Moreover, the convergence of the Neumann series (and hence the accuracy of the Born approximation) requires $\eta$ to be sufficiently small. In the case of the synthetic dataset, the Born approximation fails when $\eta_{i}\gg 1$ for any circle, thus preventing an accurate inversion of (13). In particular, a direct inversion results in severe underestimates of the true contrast regardless of the chosen regularization parameter $\gamma$, as seen in Figure 3. When weak scattering breaks down, the Born approximation also gives rise to artifacts when there is multiple scattering (see Figure 3). We seek to remediate the failures of the Born approximation in the strong scattering case through several different corrective approaches using neural networks.
Figure 3: Here we show two examples of difficulties with the simple inverse Born approximation with $\gamma=1/10$ and $\gamma=1$. Top row: The Born reconstruction for three weak ($\eta=1.2$), circular scatterers. Two of the circles are too close to be distinguished. Bottom row: The Born reconstruction of a strong ($\eta=1.8$), U-shaped resonance structure. The method produces a spurious reconstruction of the inlet, as well as incorrect contrasts nearby.
@@ 4.1 Correction strategies
We propose two methods for correcting the Born approximation: a pre-correction and a post-correction. Let $\vec{\mu}_{\rm exact}$ denote the exact nodal values for the (vectorized) contrast. There is an error $\tau_{\mu}$ associated with both the convergence of the Neumann series and its truncation (the Born approximation). We may write
$\mathcal{B}\vec{\mu}_{\rm exact}=\vec{u}_{\infty}+\tau_{\mu}\mathrel{\mathop{:}}=\vec{u}_{\infty}^{\tau_{\mu}}.$ (16)
where $\mathcal{B}$ is the stacked Born operator defined in (14), and $\vec{u}_{\infty}$ is the exact far field pattern. The factor $\tau_{\mu}$ can be thought of as the discrepancy in the Born-obtained far field compared to the true (simulated) far field. A CNN is trained to predict $\tau_{\mu}$, and this approach pre-corrects the far field data into a form suitable for the Born approximation to predict an accurate contrast. In particular
@@
provided that $\gamma$ is chosen correctly and $\vec{u}_{\infty}^{\tau_{\mu}}$ is learned appropriately. This is motivated by the approach in [14] where a neural network is used to correct data from approximate forward models used to speed up forward computations.
@@
This is the approach used in [22, 15] in a different context for correcting satellite data-based retrieval algorithms.
@@ 4.2 Training
    $\mathcal{L}_{\rm CNNB}(\beta)=\frac{1}{N_{\rm train}}\sum_{i=1}^{N_{\rm train}}\|\vec{u}^{\tau_{\mu}(i)}_{\infty}-\operatorname{CNNB}(\vec{u}_{\infty}^{(i)};\beta)\|^{2},$ (18)
    where $\rm CNNB(\>\cdot\>;\beta)$ is a CNN with weights $\beta$.
  • BCNN: We consider data pairs $\{(\vec{u}_{\infty}^{(i)},\varepsilon_{\mu}^{(i)}\}_{i=1}^{N_{\rm train}}$ and seek to minimize the loss function
    $\mathcal{L}_{\rm BCNN}(\beta^{\prime})=\frac{1}{N_{\rm train}}\sum_{i=1}^{N_{\rm train}}\|\varepsilon_{\mu}^{(i)}-\operatorname{BCNN}(\vec{u}_{\infty}^{(i)};\beta^{\prime})\|^{2},$ (19)
    where $\rm BCNN(\>\cdot\>;\beta^{\prime})$ is another CNN with weights $\beta^{\prime}$.
@@
$\mathcal{L}_{\rm CNN}(\beta^{\prime\prime})=\frac{1}{N_{\rm train}}\sum_{i=1}^{N_{\rm train}}\|\vec{\mu}^{(i)}-\rm CNN(u_{\infty}^{(i)};\beta^{\prime\prime})\|^{2},$ (20)
where $\rm CNN(\>\cdot\>,\beta^{\prime\prime})$ is a third CNN with weights $\beta^{\prime\prime}$. To aid comparison we use the same general CNN structure in all three cases, but optimize the network hyper-parameters to each case. Furthermore, we compute the basic regularized Born inverse (15) with regularization parameter $\gamma=1$ and $\gamma=0.1$ for a direct comparison of CNN results to the Born approximation.
@@
Figure 4: The neural network architecture for the three models compared in this paper. (a) The CNN model directly maps the far field to the squared refractive index. (b) The regularized Born inverse is computed to obtain a rough estimate of the squared refractive index $\eta(\textbf{x})$, which is then corrected by the BCNN predicted $\varepsilon$. (c) The CNNB model pre-corrects the far field and then applies a regularized Born inverse to obtain the squared refractive index. Note that the necessary reshaping of the far field are omitted in this diagram.
In all three approaches, the model takes as input a representation of the far-field matrix, specifically, where the real and imaginary parts are separated and stored in two distinct channels, which enables the model to process both components simultaneously. The model output is then reshaped into a two-dimensional map corresponding to the discrepancy term. This end-to-end mapping from far-field data to an image-like discrepancy allows the model to leverage spatial correlations in the data that are challenging to capture through purely analytic approaches. The structure of the three methods is visualized in Figure 4.

### v2

# A Neural Network Enhanced Born Approximation for Inverse Scattering

@@ Abstract
Time-harmonic, acoustic inverse scattering concerns the ill-posed and nonlinear problem of determining the refractive index of an inaccessible, penetrable scatterer based on far field wave scattering data. When the scattering is weak, the Born approximation provides a linearized model for recovering the shape and material properties of a scatterer. We develop two neural network algorithms–Born-CNN (BCNN) and CNN-Born (CNNB)–to correct the Born approximation when the scattering is not weak. BCNN applies a post-correction to the Born reconstruction, while CNNB pre-corrects the data. Both methods leverage the Born approximation’s excellent fidelity in weak scattering, while extending its applicability beyond its theoretical limits. CNNB particularly exhibits a strong generalization to noisy and absorbing scatterers. Based on numerical tests, our approach provides alternative data-driven methods for obtaining the refractive index, extending the utility of the Born approximation to regimes where the traditional method fails.
@@ 1 Introduction
The time-harmonic inverse scattering problem concerns the determination of the properties, such as shape or material composition, of an inaccessible object from remote measurements of wave scattering data. Problems of this type arise in, for example, seismology, medical imaging, and radar applications. Due to the significance of the applications, there has been extensive work on various algorithms for this problem (see for example [6]).
We shall study a particular inverse scattering problem: the inverse medium problem for the Helmholtz equation. In this case, it is desired to reconstruct the refractive index of a bounded penetrable scatterer from far field acoustic data. The major complication is that this inverse scattering problem is both ill-posed and nonlinear [6, Theorem 4.21, page 448]. Current approaches can be divided into two broad classes: 1) quantitative methods and 2) qualitative methods. A quantitative method attempts to reconstruct the scatterer directly. Often this involves a non-linear and regularized optimization problem that seeks a reconstruction corresponding to a far field pattern that matches measurement data. This is computationally intensive and can fail due to local minima [6]. Another quantitative approach is to assume that the scattering is weak and to use the inverse Born approximation. We shall discuss this in more detail shortly since correcting this approach is the subject of our paper. In contrast to these quantitative methods, the goal of a qualitative method such as the Linear Sampling Method [3, 6] is to approximate the support or boundary of the scatterer. Approaches of this type do not involve optimization or the solution of the forward problem. They do not require strong a priori assumptions, but cannot directly distinguish different materials in the field of view.
@@
The novelty of this paper is to suggest and test two schemes combining the Born approximation with a neural network to extend the domain of applicability of the Born approximation. In particular, we compare the resulting predictions to the standard regularized inverse Born approximation, and to a CNN trained simply to invert the data along the lines of [10]. After optimizing the networks and training on simple data generated by a few circular scatterers, we show that the so called CNNB model is remarkably stable to added noise.
Limited testing on more complex, out-of-distribution scatterers reveals that all methods improve the fidelity of the reconstructions for strong scatterers, but the inclusion of the Born approximation in the two proposed schemes produces generally better results. The study suggests that combining CNNs and the Born approximation has promise in solving inverse scattering problems and that the training phase does not need to include close copies of the scatterers.
@@ 2 The Inverse scattering problem
Figure 1: An illustration of the scattering problem. The incident field $u^{i}$ travels with direction of propagation $\varphi$ and interacts with the compactly supported scatterer given by $\eta(\textbf{x})\not=1$ shown in blue, and contained in a bounded set $\Omega$ (shown in gray). The scattered field $u^{s}$ propagates outward with direction angle $\theta$.
@@
$u^{i}(\mathbf{x},\mathbf{d})=\exp(\mathrm{i}\mkern 1.0muk\mathbf{x}\cdot\mathbf{d}),\quad\mathbf{d}=\langle\cos\varphi,\sin\varphi\rangle$ (1)
@@
$\Delta u+k^{2}\eta(\mathbf{x})u=0\text{ in }$ $\mathbb{R}^{2},$ (2)
@@
Under the conditions given above, Equations (2)-(4) have a unique solution for any $k>0$ [6]. The Forward Problem consists of solving the above linear well-posed problem given $\mathbf{d}$, $k$, and $\eta$ (see Fig. 1).
It follows from the fact that $u^{s}$ satisfies the Helmholtz equation and the radiation condition that $u^{s}$ exhibits an asymptotic expansion as an outgoing cylindrical wave for $|\mathbf{x}|$ sufficiently large:
@@
The Inverse Problem that we wish to solve is to determine $\mu(\textbf{x})$ (equivalently, $\eta(\mathbf{x})$, the square of the refractive index) given the far field pattern $u_{\infty}(\hat{\mathbf{x}},\mathbf{d})$ for all $\hat{\mathbf{x}}$ and d on $\mathbb{S}^{1}$ (in practice, only a finite set of $\hat{\mathbf{x}}$ and d are used). Here we assume data is given for a single fixed wave number $k>0$, and that the support of $\mu$ is a priori known to lie in a bounded search region $\Omega$ (see Fig. 1). This problem is non-linear and ill-posed [6].
In this study, we use synthetic scattering data generated through a standard finite element approach to approximating (2)-(4). For each incident direction d, the total field $u$ in a neighborhood of the scatterer is computed using the Netgen package [21] with 4th-order elements and a mesh-size request of one-eighth of the local wavelength of the wave. The boundary of each scatterer is fitted using isoparametric curved elements. The Sommerfeld radiation condition is handled through a radial Perfectly Matched Layer (PML) implemented by Netgen using complex stretching as discussed in [5]. To generate an approximate far field pattern, we follow [18] to map the near field to a far field pattern.
@@ 2.1 The Born approximation
$H^{1}_{loc}(\mathbb{R}^{2})=\{u:\mathbb{R}^{2}\to\mathbb{C}\>|\>u|_{B_{t}}\in H^{1}(B_{t})\>\>\forall\>\>t>0\},$ (6)
where $B_{t}=B(t,0)$ denotes a disk of radius $t$ centered at the origin. In particular, let $\Omega=B_{R}$ be the smallest such disk containing the support of $\mu$. It can be shown that if $u\in H^{1}_{loc}(\mathbb{R}^{2})$ is a solution to the scattering problem (2)-(4), then $u|_{\Omega}\in L^{2}(\Omega)$ and satisfies the Lippmann-Schwinger equation:
@@
provided that the operator norm $\|T\|_{\infty}<1$. A sufficient condition for this is [19]
@@
where ${\mathscr{B}}^{*}$ is the $L^{2}$-adjoint of ${\mathscr{B}}$, $I$ is the identity operator and $\gamma>0$ is a fixed regularization parameter determined a priori. Note that fast methods exist to compute this approximation (including the NN approach of Zhou [26] or the low rank approximation method of [25]) but we do not use them here.
@@ 3.1 Discretization of the Born approximation
$(U_{\infty})_{i,j}\approx\frac{k^{2}}{\sqrt{8\pi k}}\exp\left(\frac{\mathrm{i}\mkern 1.0mu\pi}{4}\right)h^{2}\sum_{p,q=1}^{n}\exp(\mathrm{i}\mkern 1.0muk(\textbf{d}_{j}-\hat{\textbf{x}}_{i})\cdot\textbf{y}_{p,q})M_{p,q},$ (13)
@@
where $\mathscr{B}_{\rm disc}$ encodes the (2+2)-discretization in both angular dimensions and both spatial dimensions, and $\tau_{\mu}$ is the unknown error for the Born approximation.
To utilize linear algebraic methods, we can use standard tensor unfolding methods to collapse $\mathscr{B}_{\rm disc}$ along the spatial and angular dimensions to produce the aforementioned stacked Born operator $\mathcal{B}\in\mathbb{C}^{N_{R}\times N_{S}\times N_{g}^{2}}$. We thus arrive at the following formulation:
@@
where $\vec{u}_{\infty}=\operatorname{vec}(U_{\infty})$, $\vec{\mu}=\operatorname{vec}(M)$, and $\vec{\tau}_{\mu}=\operatorname{vec}(\tau_{\mu})$ are the vectorized formats (through appropriate reshaping) of the far field, contrast, and the unknown error respectively. The discrete inverse Born approximation then predicts the nodal values of $\mu$ by ignoring the error term $\vec{\tau}_{\mu}$ and using
@@ 3.2 Data generation
Figure 2: Two realizations of the random training data. Each row displays a density plot of $\eta$ for a sample scatterer (left panel) alongside the real and imaginary components of its computed far field pattern. The far field pattern is the data for the inverse scattering problem.
For training, we want to use simple scatterers. Here we create the scatterer field $\eta(\textbf{x})$ as the union of $N_{c}$ piecewise-constant circles overlaying a homogeneous background of air, where we uniformly sample $N_{c}\in\{1,2,3\}$. The $i$th circle is assigned a constant value $\eta_{i}$ sampled from the uniform distribution $\mathcal{U}(1.1,1.8)$, with the background air set to $\eta=1$ (we shall see that $\eta\approx 1.8$ is outside the weak scattering regime). The radius and position of each circle are sampled uniformly from $\mathcal{U}(0.1,0.3)$ and $\mathcal{U}(-0.7,0.7)$, respectively. In selecting the training data, for simplicity of mesh generation, we enforce that the circles cannot overlap and must be fully contained in the search domain $[-1,1]^{2}$. We generate 20,000 samples for training and validation (under an 80-20% split) with an additional 4,000 samples for testing purposes. Visualizations of typical training samples can be seen in Figure 2.
The validity of the Born approximation is contingent upon weak scattering of the incident field. Moreover, the convergence of the Neumann series (and hence the accuracy of the Born approximation) requires $\eta$ to be sufficiently close to unity. In the case of the synthetic dataset, the Born approximation fails when $\eta_{i}\gg 1$ for any circle, thus preventing an accurate inversion of (14). In particular, a direct inversion results in severe underestimates of the true contrast regardless of the chosen regularization parameter $\gamma$.
When weak scattering breaks down, the Born approximation also gives rise to artifacts when there is multiple scattering. Examples of poor reconstructions using the Born approximation are shown in Section 5. We seek to remediate these failures in the strong scattering case through corrective approaches using neural networks.
@@ 4 Network architecture and training
Throughout the rest of the paper we shall used two norms defined on vectors with $n$ components so that if $\vec{v}\in\mathbb{C}^{n}$ then the $\ell^{p}$ norm of $\vec{v}$ is
$\|\vec{v}\|_{p}=\left(\sum_{j=1}^{n}|v_{j}|^{p}\right)^{1/p},\;p=1,2.$
For arrays (in particular the array of pixel intensities for the image), we first reshape the array into a vector and then compute the corresponding vector norm as above. By abuse of notation we shall use $\|\cdot\|_{p}$ to indicate the norm of a vector or a matrix ($p=2$ is in fact the Frobenius norm of a matrix).
@@ 4.1 Correction strategies
We propose two methods for correcting the Born approximation: a pre-correction and a post-correction. Let $\vec{\mu}_{\rm exact}$ denote the exact nodal values for the (vectorized) contrast. There is an error $\vec{\tau}_{\mu}$ associated with both the convergence of the Neumann series and its truncation (the Born approximation). We may write
$\mathcal{B}\vec{\mu}_{\rm exact}=\vec{u}_{\infty}+\vec{\tau}_{\mu}\mathrel{\mathop{:}}=\vec{u}_{\infty}^{\tau_{\mu}}.$ (17)
where $\mathcal{B}$ is the stacked Born operator defined in (15), and $\vec{u}_{\infty}$ is the exact far field pattern. The factor $\vec{\tau}_{\mu}$ can be thought of as the discrepancy in the Born-obtained far field compared to the true (simulated) far field. A CNN is trained to predict $\vec{\tau}_{\mu}$, and this approach pre-corrects the far field data into a form suitable for the Born approximation to predict an accurate contrast. In particular
@@
provided that $\gamma$ is chosen correctly and $\vec{u}_{\infty}^{\tau_{\mu}}$ is learned appropriately. This is motivated by the approach in [15], where a neural network is employed to compensate for modeling errors introduced by approximate forward models.
@@
This is the approach used in, for example, [23, 16] in a different context for correcting satellite data-based retrieval algorithms.
@@ 4.2 Training
    $\mathcal{L}_{\rm CNNB}(\beta)=\frac{1}{N_{\rm train}}\sum_{i=1}^{N_{\rm train}}\|\vec{u}^{\tau_{\mu}(i)}_{\infty}-\operatorname{CNNB}(\vec{u}_{\infty}^{(i)};\beta)\|_{2}^{2},$ (19)
    where $\rm CNNB(\>\cdot\>;\beta)$ is a CNN with weights and biases collected in the vector $\beta$.
  • BCNN: We consider data pairs $\{(\vec{u}_{\infty}^{(i)},\varepsilon_{\mu}^{(i))}\}_{i=1}^{N_{\rm train}}$ and seek to minimize the loss function
    $\mathcal{L}_{\rm BCNN}(\beta^{\prime})=\frac{1}{N_{\rm train}}\sum_{i=1}^{N_{\rm train}}\|\varepsilon_{\mu}^{(i)}-\operatorname{BCNN}(\vec{u}_{\infty}^{(i)};\beta^{\prime})\|_{2}^{2},$ (20)
    where $\rm BCNN(\>\cdot\>;\beta^{\prime})$ is another CNN weights and biases collected in the vector $\beta^{\prime}$.
@@
$\mathcal{L}_{\rm CNN}(\beta^{\prime\prime})=\frac{1}{N_{\rm train}}\sum_{i=1}^{N_{\rm train}}\|\vec{\mu}^{(i)}-{\rm CNN}(u_{\infty}^{(i)};\beta^{\prime\prime})\|_{2}^{2},$ (21)
where $\rm CNN(\>\cdot\>,\beta^{\prime\prime})$ is a third CNN with weights and biases $\beta^{\prime\prime}$.
To aid comparison we use the same general CNN structure in all three cases, but optimize the network hyper-parameters to each case. Furthermore, we compute the basic regularized Born inverse (16) with regularization parameter $\gamma=1$ and $\gamma=0.1$ for a direct comparison of CNN results to the Born approximation.
@@
Figure 3: The neural network architecture for the three models compared in this paper. (a) The CNN model directly maps the far field to the squared refractive index. (b) The regularized Born inverse is computed to obtain a rough estimate of the squared refractive index $\eta(\textbf{x})$, which is then corrected by the BCNN predicted $\varepsilon$. (c) The CNNB model pre-corrects the far field and then applies a regularized Born inverse to obtain the squared refractive index. Note that the necessary reshaping of the far field are omitted in this diagram, as well as the map $\mu\mapsto\mu+1=\eta$.
In all three approaches, the model takes as input a representation of the far field matrix, specifically, where the real and imaginary parts are separated and stored in two distinct channels, which enables the model to process both components simultaneously. The model output is then reshaped into a two-dimensional map corresponding to the discrepancy term. This end-to-end mapping from far field data to an image-like discrepancy allows the model to leverage spatial correlations in the data that are challenging to capture through purely analytic approaches. The structure of the three methods is visualized in Figure 3.